### v1

# Towards efficient quantum algorithms for diffusion probability models

Towards efficient quantum algorithms for diffusion probability models
Yunfei Wang Equal contribution Joint Center for Quantum Information and Computer Science, NIST/University of Maryland, College Park, MD 20742, USA Maryland Center for Fundamental Physics, Department of Physics, University of Maryland, College Park, MD 20742, USA    Ruoxi Jiang Equal contribution Department of Computer Science, University of Chicago Chicago, IL 60637    Yingda Fan Department of Computer Science, The University of Pittsburgh, Pittsburgh, PA 15260, USA    Xiaowei Jia Department of Computer Science, The University of Pittsburgh, Pittsburgh, PA 15260, USA    Jens Eisert jense@zedat.fu-berlin.de Dahlem Center for Complex Quantum Systems, Freie Universität Berlin, 14195 Berlin, Germany    Junyu Liu junyuliu@pitt.edu Department of Computer Science, The University of Pittsburgh, Pittsburgh, PA 15260, USA    Jin-Peng Liu liujinpeng@tsinghua.edu.cn Yau Mathematical Sciences Center and Department of Mathematics, Tsinghua University, Beijing 100084, China Yanqi Lake Beijing Institute of Mathematical Sciences and Applications, Beijing 100407, China
A diffusion probabilistic model (DPM) is a generative model renowned for its ability to produce high-quality outputs in tasks such as image and audio generation. However, training DPMs on large, high-dimensional datasets—such as high-resolution images or audio—incurs significant computational, energy, and hardware costs. In this work, we introduce efficient quantum algorithms for implementing DPMs through various quantum ODE solvers. These algorithms highlight the potential of quantum Carleman linearization for diverse mathematical structures, leveraging state-of-the-art quantum linear system solvers (QLSS) or linear combination of Hamiltonian simulations (LCHS). Specifically, we focus on two approaches: DPM-solver-$k$ which employs exact $k$-th order derivatives to compute a polynomial approximation of $\epsilon_{\theta}(x_{\lambda},\lambda)$; and UniPC which uses finite difference of $\epsilon_{\theta}(x_{\lambda},\lambda)$ at different points $(x_{s_{m}},\lambda_{s_{m}})$ to approximate higher-order derivatives. As such, this work represents one of the most direct and pragmatic applications of quantum algorithms to large-scale machine learning models, presumably talking substantial steps towards demonstrating the practical utility of quantum computing.
I Introduction
Diffusion models, in the context of machine learning and, more specifically, deep learning, represent a class of generative models that have garnered significant attention for their novel approach to modeling complex data distributions [1, 2, 3, 3, 4]. Unlike traditional models of machine learning, which often rely on explicit likelihood estimation or adversarial training, diffusion models generate data through a gradual, iterative process inspired by the physical concept of diffusion. They work by initially introducing noise a data points and then learning how to reverse this process, effectively denoising the data step-by-step. This framework allows them to generate high-quality samples, often rivaling or surpassing the performance of other generative models, such as generative adversarial networks (GANs) [5] and variational autoencoders (VAEs) [6]. The appeal of diffusion models lies in their stability during training and their ability to model intricate, high-dimensional data distributions without requiring complex adversarial setups or delicate balancing of training dynamics. Diffusion probabilistic models (DPM), as we call them, are families of generative model renowned for their ability to produce high-quality out- puts in complex tasks such as image and audio generation.
Turning to quantum approaches, quantum computing is widely regarded as one of the most promising alternatives to von Neumann architectures in the post-Moore era [7]. By leveraging unique quantum features such as superposition and entanglement, quantum computing is anticipated to surpass classical counterparts in specific tasks. These include factoring [8], database search [9], simulating complex quantum systems [10], sampling [11], and solving problems in linear algebra [12, 13]. Although we are currently still in what can be called the noisy intermediate-scale quantum (NISQ) era, the concept of a megaquop machine offers a vision of a future where quantum computing achieves remarkable reliability and scalability. Such a machine, with an error rate per logical gate on the order of $10^{-6}$, would surpass the limitations of classical, NISQ, or analog quantum devices. It would have the capacity to execute quantum circuits involving around $100$ logical qubits and a circuit depth of approximately $10,000$—capabilities far beyond the reach of current technology.
Figure 1: Sketch of classical (1.) diffusion probability models (DPM) and the quantum algorithms proposed here for quantum analogs (2.-4.). We pursue two approaches: This is on the one hand Carlemann linearization for DPM solvers (2.), on the other Carleman linearization for UniPC (3.). For both approaches, quantum ODE solvers are instrumental in the step before measurement. QLSS stands for quantum linear system solvers, LCHS for linear combination of Hamiltonian simulation.
In the light of these observations, it is not surprising that recent years have seen a surge of interest in exploring how quantum algorithms might enhance machine learning tasks, whether in terms of sample complexity, computational complexity, or generalization [14, 15, 16, 12]. Indeed, This interest is well motivated by the fact that, for certain structured models, quantum algorithms are proven to vastly outperform their classical counterparts. In fact, rigorous separations have been demonstrated for specific PAC learning problems [17, 18, 19]. Even constant-depth quantum circuits have been shown to offer advantages over suitably constrained classical circuits [20]. These insights provide strong motivation for the rapidly evolving field of quantum machine learning and give impetus for the thought that quantum computers may well offer computational advantages.
That said, quantum computers excel at tackling highly structured problems, while many common machine learning challenges involve highly unstructured data. In the light of this, it remains to an extent unclear what practical advantages quantum computers might offer for addressing machine learning tasks. Furthermore, it is still uncertain whether proving separations will be the most effective or even the best approach for advancing the field [21]. For variational approaches [16], while being exciting and promising, no rigorous guarantees and separations of quantum over classical approaches are known yet. To date, it seems fair to say that are lacking a framework for devising end-to-end applications of quantum computers for industrially relevant machine learning tasks [22].
In this work, we take out-of-the-box steps to establish a fresh avenue towards showing the utility of quantum computers in machine learning tasks. Concretely, we present novel and efficient quantum algorithms designed for implementing differential path models (DPMs) through a variety of quantum ordinary differential equation (ODE) solvers. Our algorithms showcase the potential of quantum Carleman linearization in tackling a broad range of mathematical structures and problems, leveraging some of the latest advancements in quantum linear system solvers (QLSS) or linear combination of Hamiltonian simulation (LCHS). Specifically, we explore two distinctly different but related approaches: the DPM-solver-$k$, which utilizes exact $k$-th order derivatives to compute a polynomial approximation of $\epsilon_{\theta}(x_{\lambda},\lambda)$, and UniPC, which approximates higher-order derivatives by employing finite differences of $\epsilon_{\theta}(x_{\lambda},\lambda)$ at multiple points $(x_{s_{m}},\lambda_{s_{m}})$. Furthermore, this work represents one of the most direct and pragmatic applications of quantum algorithms to large-scale machine learning models, highlighting a key milestone in demonstrating the tangible benefits and practical utility of quantum computing in real-world tasks. We provide both rigorous statements and numerical results. While this result may not yet be a full end-to-end application, by bridging the gap between theoretical quantum algorithms and their application to complex computational models, we take a significant step, so we think, toward realizing the power of quantum technologies for solving real-world problems in machine learning and beyond.
II Results
In this section, we lay out our specific results, both concerning rigorous statements that heuristic numerical statements that are meant to provide further evidence for the functioning of the approach. Substantial more detail on all steps will be presented in the supplemental material. We will now explain what steps we take towards establishing a fresh direction for demonstrating the utility of quantum computers in machine learning tasks. Specifically, we present novel and efficient quantum algorithms for implementing diffusion probabilistic models (DPMs) using a variety of ordinary differential equation (ODE) solvers. These algorithms leverage recent advancements in quantum ODE solvers and highlight the potential of quantum Carleman linearization for addressing a broad range of mathematical structures and challenges. Our exploration focuses on two distinct approaches.
  • DPM-solver-$k$: Utilizes exact $k$-th order derivatives to compute a polynomial approximation of $\epsilon_{\theta}(x_{\lambda},\lambda)$.
  • UniPC-$p$: Approximates higher-order derivatives by employing finite differences of $\epsilon_{\theta}(x_{\lambda},\lambda)$ at different points $(x_{s_{m}},\lambda_{s_{m}})$.
These algorithms underscore the potential of quantum Carleman linearization in addressing diverse mathematical structures. Moreover, as we advance toward the era of megaquop machines, this work stands out as one of the most direct and practical applications of quantum algorithms to large-scale machine learning models. It marks, so we hope, a significant milestone in showcasing the tangible utility of quantum computing in real-world applications.
II.1 Embedding classical neural networks
In the center of DPMs in the classical realm are ordinary differential equations capturing diffusion processes. On the highest level, they are defined by a forward process, the reverse or backward process, and the sampling procedure. In this work, the core aim is to establish quantum algorithms for quantum analogs of such models. Since quantum mechanics is intrinsically linear, much of the work will circle around embedding the problem in the appropriate fashion. In this section, we present a brief heuristic overview of how to solve the diffusion ODE
$\frac{\mathrm{d}x_{t}}{\mathrm{d}t}=f(t)x_{t}+\frac{g^{2}(t)}{2\sigma_{t}}\epsilon_{\theta}(x_{t},t),\qquad x_{T}\sim q_{T}(x_{T}),$ (1)
leveraging an instance of a quantum ODE solver based on both the DPM-solver [23, 24] and UniPC [25].
II.1.1 Carleman linearization for DPM-solvers
We here outline the core ideas of the linearization and embedding – details can be found in the supplemental material, also shown in Fig. 1. The DPM-solver series are currently the most widely used solvers for diffusion ODEs. Given an initial value $x_{s}$, the solution $x_{t}$ of the diffusion ODE w.r.t. the noise prediction model Eq. (1) for $t\in(0,s)$ is given by
$\frac{x_{t}}{\alpha_{t}}=\frac{x_{s}}{\alpha_{s}}-\int_{\lambda_{s}}^{\lambda_{t}}e^{-\lambda}\epsilon_{\theta}(x_{\lambda},\lambda)\mathrm{d}\lambda.$ (2)
In this expression $\lambda\coloneqq\log(\alpha_{t}/\sigma_{t})$ is what is called the log-SNR variable. For a given a set of time steps $\{t_{i}\}_{i=0}^{M}$ with $t_{0}=T$ and $t_{M}=0$, with $h_{i}=\lambda_{t_{i}}-\lambda_{t_{i-1}}$, one makes use of the $(k-1)$-st-order Taylor approximation in order to develop what is called the DPM-solver-$k$ in Ref. [23]. This actually takes the form
$\begin{split}x_{t_{i}}&=\frac{\alpha_{t_{i}}}{\alpha_{t_{i-1}}}x_{t_{i-1}}-\alpha_{t_{i}}\sum_{n=0}^{k-1}\epsilon^{(n)}_{\theta}(x_{\lambda_{t_{i-1}}},\lambda_{t_{i-1}})\\ &\times\int_{\lambda_{s}}^{\lambda_{t}}e^{-\lambda}\frac{(\lambda-\lambda_{t_{i-1}})^{n}}{n!}\mathrm{d}\lambda+O(h_{i}^{k+1}),\end{split}$ (3)
where $\epsilon^{(n)}_{\theta}(x_{\lambda_{t_{i-1}}},\lambda_{t_{i-1}})$ is the $n$-th derivative. The integral in the final line can be computed analytically in this case. In the same way, a polynomial expansion on the data prediction model $x_{\theta}(x_{\lambda},\lambda)$ gives rise to what is called the DPM-solver++ [24]. However, as mentioned before, the intrinsic linearity of quantum mechanics renders a direct implementation of non-linear ODE schemes inapplicable, so that one has to resort to an instance of Carleman or Koopman linearization and embedding of the diffusion ODE. In what follows, we discuss the logic and structure of the proposed quantum algorithm.
  • In a first step, we express $\epsilon_{\theta}(x_{\lambda},\lambda)$ and its higher-order derivatives as a polynomial expansion in $x_{\lambda}$. Building on this, we derive a forward iteration $x_{t_{i}}=\operatorname{poly}(x_{t_{i-1}})$.
  • For this still being a non-linear expression, we employ Carleman linearization to transform it into a linear system, so that a quantum algorithm can be devised.
  • For this, we define $y_{j}(\lambda)\approx x^{j}_{\lambda}$ for $j=1,\dots,N$. As the Carleman linearization transforms a non-linear differential equation into an infinite-dimensional system of linear differential equations, we truncate it at step $N$, with the initial condition $y_{j}(\lambda_{t_{i-1}})=x^{j}_{\lambda_{t_{i-1}}}$.
  • Setting $Y=(y_{1},\ldots,y_{N})$, we then derive a linear iteration of the form
    $\begin{split}\delta\hat{Y}&=\hat{Y}(t_{i})-\hat{Y}(t_{i-1})=A\hat{Y}(t_{i-1})+b\leavevmode\nobreak\ ,\\ &\text{ for }\quad\hat{y}_{1}(t_{M})=\hat{y}_{1}(0)=x_{\mathrm{in}}(0)\leavevmode\nobreak\ ,\end{split}$ (4)
    based on Eq. (3), introducing a matrix $A$, which is the quantum Carleman matrix (QCM) associated with the DPM-solver.
Eq. (4) embodies $M+1$ iterations in total ranging from $t_{0}=T$ to $t_{M}=0$. Here, $X_{\text{in}}=X(0)$ is specified by the initial vector. This expression, distinctly departing from that of Ref. [26], has been designed to reconstruct the original data $x_{0}$ from $x_{t}$ through a backward denoising diffusion process, contrasting with the forward progression typically seen in a (stochastic) gradient descent algorithm during the training process.
II.1.2 Carleman linearization for UniPC
Complementing the above methodology, in this subsection, we design a reading of a quantum Carleman algorithm suitable for the known alternative technique to solve the diffusion ODE referred to as UniPC. This includes both the predictor (UniP-p) and the corrector (UniC-p), defined as
$\begin{split}x_{s_{p}}=\frac{\alpha_{s_{p}}}{\alpha_{s_{0}}}x_{s_{0}}&-\sigma_{s_{p}}(e^{h_{i}}-1)\epsilon_{\theta}(x_{s_{0}},\lambda_{s_{0}})\\ &-\sigma_{s_{p}}B(h_{i})\sum_{m=1}^{p-1}\frac{a_{m}}{r_{m}}D_{m},\end{split}$ (5)
and as
$\begin{split}x^{c}_{s_{p}}=\frac{\alpha_{s_{p}}}{\alpha_{s_{0}}}x^{c}_{s_{0}}&-\sigma_{s_{p}}(e^{h_{i}}-1)\epsilon_{\theta}(x_{s_{0}},\lambda_{s_{0}})\\ &-\sigma_{s_{p}}B(h_{i})\sum_{m=1}^{p}\frac{a_{m}}{r_{m}}D_{m}.\end{split}$ (6)
Here, we have encountered the interpolation time steps $s_{m}=h_{i}r_{m}+\lambda_{t_{i-1}}$, where $m=1,2,\dots,p$, and the finite difference $D_{m}=\epsilon_{\theta}(x_{s_{m}},\lambda_{s_{m}})-\epsilon_{\theta}(x_{s_{0}},\lambda_{s_{0}})$. Note that in this expression $s_{0}=\lambda_{t_{i-1}}$, and $s_{p}=\lambda_{t_{i}}$. Again, we outline the logic and structure of the proposed quantum algorithm.
  • In a first step, as before, we rewrite the denoising function $\epsilon_{\theta}(x_{\lambda},\lambda)$ in terms of a polynomial expansion in $x_{\lambda}$ up to order $J$, with the coefficients $a_{j}$ contains $\lambda$.
  • In a next step, we can express $\epsilon_{\theta}(x_{\lambda},\lambda)$ at different points as a polynomial of $x_{\lambda}$, and derive a multi-step forward iteration for UniP given by $x_{s_{p}}=\operatorname{poly}(x_{s_{0}})+\operatorname{poly}(x_{s_{1}})+\cdots+\operatorname{poly}(x_{s_{p-1}})$. For UniC, this is $x^{c}_{s_{p}}=x^{c}_{s_{0}}+\operatorname{poly}(x_{s_{0}})+\operatorname{poly}(x_{s_{1}})+\cdots+\operatorname{poly}(x_{s_{p}})$.
  • Similarly as before, we use Carleman linearization in this step and truncate the system at $N$. We denote $y_{j}(\lambda)\approx x^{j}_{\lambda}(\lambda)$, $z_{j}(\lambda)\approx(x^{c})^{j}_{\lambda}(\lambda)$ for $j=1,\dots,N$, with the initial condition $y_{1}(s_{0})=x_{\lambda}(s_{0})$, $z_{1}(s_{0})=(x^{c})_{\lambda}(s_{0})$.
  • In terms of $Y=(y_{1},\ldots,y_{N})$, and $Z=(z_{1},\ldots,z_{N})$, we can build on Eq. (5) and Eq. (6) to derive a multi-step linear iteration (predictor) expression given by
    $\begin{split}\delta Y(t_{i})=\leavevmode\nobreak\ &A^{(0)}Y^{(0)}(t_{i})+b^{(0)}\\ &+A^{(1)}Y^{(1)}(t_{i})+b^{(1)}\\ &+\ldots+A^{(p-1)}Y^{(p-1)}(t_{i})+b^{(p-1)}\leavevmode\nobreak\ ,\end{split}$ (7)
    with Carleman matrices $A^{(0)},\dots,A^{(p-1)}$ for all interpolation steps. The corrector, reflecting the other linear iteration takes the form
    $\begin{split}\delta Z(t_{i})=BZ(t_{i})&+A^{(0)_{c}}Y^{(0)}(t_{i})+b^{(0)_{c}}\\ &+A^{(1)_{c}}Y^{(1)}(t_{i})+b^{(1)_{c}}\\ &+\ldots+A^{(p)_{c}}Y^{(p)}(t_{i})+b^{(p)_{c}}\leavevmode\nobreak\ ,\end{split}$ (8)
    with Carleman matrices $B,A^{(0)_{c}},\dots,A^{(p)_{c}}$.
It is important to stress that the Carleman linearization approach introduced here represents a novel Carleman predictor-corrector scheme. This approach differs from both the DPM-solver framework described in the previous section and the approach taken in Ref. [26]. A defining feature of the quantum Carleman algorithm for UniPC is the generation of a unique quantum Carleman matrix (QCM) for each interpolation time step $s_{m}$ that we think is also interesting as an innovative step in its own right.
Figure 2: Results of numerical experiments. 1. Spectrum dynamics of the reverse diffusion process on ImageNet-100. Using the DPM-Solver [23], we show the eigenvalue trajectories of the Jacobian $\mathbf{J}(t)$ for dimensions $d=1024,4096$, and $16384$, each representing image generation at a different resolution, ranging from coarse to fine-grained. Across all three cases, the eigenvalues start positive and progressively develop a larger spectral gap with extremely positive values, highlighting the system’s consistent dissipative behavior. 2. Dissipative measure of the reverse diffusion process on ImageNet-100. The shaded area represents the [25th, 75th] percentile of the statistics. Across three models with $d=$ 1024, 4096, and 16384, $P(t)$ consistently decreases with $t$.
II.2 End-to-end quantum ODE solver
With the foundational elements established, there are two distinct approaches to solving the resulting linear ODEs. The first approach leverages the advanced capabilities of the quantum linear system solver (QLSS), while the second utilizes the linear combination of Hamiltonian simulation (LCHS) algorithm. We will begin by exploring the application of QLSS and subsequently discuss the LCHS method.
To make use of QLSS, we need to massage the iteration ODEs into a single vector equation, as has been done in Ref. [26]. We look at the DPM-solver first, the entire iteration scheme established in Eq. (4) can be captured in a vector equation
$M\bm{Y}=\bm{b}\leavevmode\nobreak$ (9)
that collects $Y(t_{i})$ at all time steps in order to form a single big vector $\bm{Y}=(Y(t_{0}),Y(t_{1}),\dots,Y(t_{M}))$. Similarly, the UniPC system can also be captured in a single vector equation as
$\bm{\mathsf{M}^{c}}\bm{\mathsf{Z}}=\bm{\mathsf{b}^{c}}\leavevmode\nobreak\ ,$ (10)
where $\bm{\mathsf{Z}}$ collects all of the $Z(t_{i})$. The explicit forms of Eq. (9) and Eq. (10) are provided in the supplementary material.
We are now in the position to invoke and discuss an instance of a quantum linear system solver (QLSS) to efficiently solve the resulting differential equations. QLSS have technically advanced significantly since the introduction of the original Harrow-Hassidim-Lloyd (HHL) algorithm proposed in 2008, which has achieved a complexity of $\mathcal{O}(\operatorname{poly}(\kappa,1/\epsilon))$ for solving sparse linear systems, where $\kappa>0$ is the condition number and $\epsilon>0$ is the desired accuracy [13]. Subsequent developments, including Ambainis’ variable-time amplitude amplification (VTAA) and linear combination of quantum walks, have improved the scaling to $\mathcal{O}(\kappa\log(1/\epsilon))$ [27, 28]. These algorithms utilize matrix ($\mathcal{O}_{M}$) and state preparation ($\mathcal{O}_{\bm{b}}$) oracles, with $\mathcal{O}_{\bm{b}}$ often incurring lower costs in practice [29]. Further recently introduced techniques, such as kernel reflection and adiabatic evolution, have further optimized the query complexities, achieving near-optimal bounds for both $\mathcal{O}_{M}$ and $\mathcal{O}_{\bm{b}}$ [30, 31]. These advancements are particularly relevant for sparse systems like those arising in the Carleman method, where the matrix sparsity $s_{M}$ and the success probability $p_{\text{succ}}$ significantly influence the overall complexity [32].
Going further, and building on the theorems and quantum algorithms discussed and introduced in Refs. [26, 30, 31, 29], we derive the formal version of Theorem 1 and Theorem 2 presented in the next section, which establish the gate complexity bounds for the quantum Carleman algorithm (for the formal version of these theorems, we refer to the supplementary material). These results integrate advancements in quantum linear system solvers, particularly leveraging the state-of-the-art methods for efficient matrix block encoding, initial state preparation, and sparsity-aware optimization, ensuring a comprehensive understanding of the computational costs associated with solving such systems.
Besides the QLSS approach, the linear combination of Hamiltonian simulation (LCHS) strategy [33, 34] provides an alternative method for solving the linear ODEs in Eqs. (9) and (10) by representing the solution to the homogeneous equation as a linear combination of Hamiltonian simulations. When a source term (e.g., $\bm{\mathsf{b}^{c}}$ in Eq. (10) and the right-hand side of Eq. (9)) is present, Duhamel’s principle can be applied [33]. LCHS achieves gate efficiency by combining linear combinations of unitaries (LCU) [35] with existing Hamiltonian simulation techniques. The method involves truncating the infinite integral to a finite interval $[-K,K]$ and discretizing it using the $M+1$ grid points. Compared to the standard QLSS approach, LCHS directly implements the time evolution operator, significantly reducing the number of state preparation oracles required.
At last, we address briefly the challenges of uploading and downloading data for our quantum algorithms, as this aspect is sometimes underappreciated. By leveraging sparsity in both data representation and training, our framework avoids the need for quantum random access memory (QRAM) [26], thus enhancing scalability and practicality while minimizing classical-to-quantum communication overhead [36, 37, 38, 39]. The downloading problem is actually more challenging than uploading, as it involves performing state tomography [40] on the resulting quantum states from our quantum algorithms [41, 42, 12]. We focus on tomographic recovery in sparse training on sparse vectors and QCM matrices from the DPM architecture, which is similar to the structure presented in Ref. [26]. The recovery process identifies $r$-sparse computational basis vectors and reconstructs the state using a modified classical shadow estimation scheme with $n$-qubit Clifford circuits [42, 43, 40], requiring $\mathcal{O}(r\log r)$ measurements only.
II.3 Theorems
The above scheme give rise to practical quantum algorithms for a fault tolerant quantum computer. In this section, we will lay out the informally formulated main theorems that are established in this work. Details can be found in the supplementary material.
Theorem 1 (Informal complexity of the DPM-solver).
For a DPM-solver with a denoising function expanded up to the $J$-th order Taylor approximation and truncated at $N$, the system of linear equations $M\bm{Y}=\bm{b}$ can be solved using a quantum algorithm. If $M$ is sparse and block-encoded efficiently, the quantum state proportional to the solution vector can be prepared with query complexity scaling as
$\mathcal{O}(J\kappa\cdot\mathrm{polylog}(N/\epsilon))\leavevmode\nobreak\ ,$ (11)
where $\kappa>0$ is an upper bound on the condition number of $M$. The algorithm achieves a success probability $p_{\text{succ}}$ greater than $1/2$ and accuracy $\epsilon>0$.
Assuming the output weight vectors are $r$-sparse with $m\coloneqq\log_{2}(N)$, the tomographic cost for transferring quantum states to classical devices is $\mathcal{O}(m^{2}r^{3}/\epsilon^{2})$, ignoring logarithmic factors. The algorithm excludes the state preparation cost, which remains efficient for sparse training.
The next statement summarizes the sample complexity of the second approach taken.
Theorem 2 (Informal complexity of the UniPC-$p$ framework).
For the UniPC-$p$ framework used in fast sampling of DPMs, a system of linear equations $\bm{\mathsf{M}^{c}}\bm{\mathsf{Z}}=\bm{\mathsf{b}^{c}}$ can be formulated and solved using a quantum algorithm. If $\bm{\mathsf{M}^{c}}$ is sparse and efficiently block-encoded, the quantum state proportional to the solution can be prepared with query complexity
$\mathcal{O}(pJ\kappa\cdot\mathrm{polylog}(N/\epsilon))\leavevmode\nobreak\ ,$ (12)
where $\kappa>0$ bounds the condition number of $\bm{\mathsf{M}^{c}}$. The algorithm achieves a success probability $p_{\text{succ}}$ greater than $1/2$ and accuracy $\epsilon>0$.
Assuming the output weight vectors are $r$-sparse with $m\coloneqq\log_{2}(N)$, the tomographic cost for transferring quantum states to classical devices is $\mathcal{O}(m^{2}r^{3}/\epsilon^{2})$, ignoring logarithmic factors. The algorithm excludes the state preparation cost, which remains efficient for sparse training.
These statements ensure that the quantum algorithms are efficiently implementable on a fault tolerant quantum computer.
II.4 Numerical analysis and experiments
Above, we present present performance guarantees of the quantum algorithms introduced. It is a core contribution of this work, however, to also make the point that the approach taken is reasonably applicable to real-world data. This is a valuable contribution, as the proven separations for machine learning problems are for highly structured and artificial data [17, 18, 19]. In this section, we perform extensive numerical experiments to show that for real-world datasets, the conditions of the above theorems are commonly satisfied for natural data and that the processes considered are sufficiently diffusive in a precise sense.
In order to do so, we employ U-ViT [44] operating in the latent space of a pre-trained AutoEncoder from Stable diffusion [45]. Working on the ImageNet-100 dataset [46], the AutoEncoder compresses input images into a lower-dimensional latent space, reducing computational overhead while preserving perceptual quality. The forward diffusion process follows the variance-preserving schedule [3]. We train the U-ViT-M for 200 epochs using the AdamW optimizer. During sampling, we employ the DPM-Solver-fast, a mix of single-order solvers [23]. Specifically, it uses 17 pre-defined saved time steps to reverse the diffusion trajectory, with each step involving intermediate substeps to approximate the continuous ODE solution, resulting in 50 number of function evaluations (NFEs).
To evaluate the dissipative property of the reverse diffusion process, we explore whether trajectories contract in phase space over time. This involves analyzing the system’s spectral properties by computing the Jacobian of its drift term as
$\bm{J}(t)=f(t)+\frac{g^{2}(t)}{2\sigma_{t}}\frac{\partial\epsilon_{\theta}(x_{t},t)}{\partial x_{t}},$ (13)
where we further compute the eigenvalues $\lambda_{i},i=1,\dots,d$, of its symmetric form $\bm{J}(t)+\bm{J}(t)^{\dagger}$. We visualize the spectral dynamics of the sampling process for well-trained diffusion models with different dimensionality. Specifically, we project images with resolutions $128\times 128$, $256\times 256$, and $512\times 512$ into the latent space of $16\times 16\times 4$, $32\times 32\times 4$, and $64\times 64\times 4$. In the light of the above rigorous statements, a quantum advantage is plausible if dissipative modes exist which are characterized by the positivity of the Hessian eigenvalues in Ref. [26]. For the reverse-time diffusion, this amounts to testing the positivity of the eigenvalues of the Jacobian. As shown in Fig. 2 1., the sampling processes at all three resolutions exhibit a consistent pattern: They begin with all positive eigenvalues and gradually transition to a state where the spectral gap increases significantly, with some eigenvalues becoming extremely positive. This growing spectral gap suggests stronger dissipativity in the system, where certain directions in the phase space contract more rapidly than others. To further verify the trend of the eigenvalues, we define $P(t)$ as a function of normalized eigenvalues as
$P(t)\coloneqq\frac{1}{d}\sum_{i=1}^{d}\Pi_{t^{\prime}=T}^{t}(1-a_{i}(t^{\prime})),$ (14)
where $a_{i}(t)\coloneqq{\lambda_{i}(t)}/{\max_{i,t}|\lambda_{i}(t)|}$. By construction, $P(t)$ grows unbounded when negative eigenvalues dominate and decays towards zero when positive eigenvalues prevail. As shown in Fig. 2 2., across different dimensionalities of the state variables, the inference process follows a consistent pattern: $P(t)$ steadily decreases as $t$ decreases, implying the emergence of large eigenvalues along the generative process, reinforcing the existence of dissipative modes, as it is required in the assumptions of the theorems.
III Discussion
In this work, we have taken steps towards identifying practical applications of quantum computers in realistic machine learning tasks, by formulating meaningful analogs of diffusion probabilistic models that constitute an increasingly important family of classical generative models, producing high quality outputs. We build on and further develop tools of quantum solvers for ordinary differential equations to make progress in presenting quantum algorithms for quantum analogs of DPMs. Our work showcases the potential of quantum Carleman linearization for a wide range of mathematical structures, utilizing cutting-edge quantum linear system solvers and linear combination of Hamiltonian simulations. We explain the respective features of the two approaches taken. For our rigorous results, we present performance guarantees, while we offer also results of numerical experiments to provide further evidence for the functioning of the approach.
At a broader level, we aim for our work to help address a key bottleneck in exploring the potential of quantum computing in machine learning. While quantum computers have demonstrated advantages for certain highly structured problems, there is a prevailing view that they may offer limited benefits for tasks in machine learning, which typically involve less structured data and more noise [21]. This work, suggests, however, that there may be significant potential for quantum computers in performing inference tasks or generative tasks of this nature. Considering that inference costs might be more than training costs in many machine learning applications, our proposed approach could pave the way for a practical example of quantum advantage in machine learning industry, while also providing a pathway to making quantum computers more useful for solving industrially relevant problems.
Acknowledgements
J. E. has been supported by the BMBF (Hybrid++, DAQC, QSolid), the QuantERA (HQCC), the BMWK (EniQmA), the DFG (CRC 183), the Quantum Flagship (PasQuans2, Millenion), the Munich Quantum Valley, Berlin Quantum, and the European Research Council (DebuQC). J.-P. L. acknowledges support from Tsinghua University and Beijing Institute of Mathematical Sciences and Applications.
See pages 1, of appendixDPM.pdf See pages 0, of appendixDPM.pdf

### v2

@@
-Towards efficient quantum algorithms for diffusion probability models
-Towards efficient quantum algorithms for diffusion probability models
-Yunfei Wang Equal contribution Joint Center for Quantum Information and Computer Science, NIST/University of Maryland, College Park, MD 20742, USA Maryland Center for Fundamental Physics, Department of Physics, University of Maryland, College Park, MD 20742, USA    Ruoxi Jiang Equal contribution Department of Computer Science, University of Chicago Chicago, IL 60637    Yingda Fan Department of Computer Science, The University of Pittsburgh, Pittsburgh, PA 15260, USA    Xiaowei Jia Department of Computer Science, The University of Pittsburgh, Pittsburgh, PA 15260, USA    Jens Eisert jense@zedat.fu-berlin.de Dahlem Center for Complex Quantum Systems, Freie Universität Berlin, 14195 Berlin, Germany    Junyu Liu junyuliu@pitt.edu Department of Computer Science, The University of Pittsburgh, Pittsburgh, PA 15260, USA    Jin-Peng Liu liujinpeng@tsinghua.edu.cn Yau Mathematical Sciences Center and Department of Mathematics, Tsinghua University, Beijing 100084, China Yanqi Lake Beijing Institute of Mathematical Sciences and Applications, Beijing 100407, China
+Towards efficient quantum algorithms for diffusion probabilistic models
+Abstract
 A diffusion probabilistic model (DPM) is a generative model renowned for its ability to produce high-quality outputs in tasks such as image and audio generation.
-However, training DPMs on large, high-dimensional datasets—such as high-resolution images or audio—incurs significant computational, energy, and hardware costs.
+However, training DPMs on large, high-dimensional datasets such as high-resolution images or audio incurs significant computational, energy, and hardware costs.
 In this work, we introduce efficient quantum algorithms for implementing DPMs through various quantum ODE solvers.
@@
 Specifically, we focus on two approaches: DPM-solver-$k$ which employs exact $k$-th order derivatives to compute a polynomial approximation of $\epsilon_{\theta}(x_{\lambda},\lambda)$; and UniPC which uses finite difference of $\epsilon_{\theta}(x_{\lambda},\lambda)$ at different points $(x_{s_{m}},\lambda_{s_{m}})$ to approximate higher-order derivatives.
-As such, this work represents one of the most direct and pragmatic applications of quantum algorithms to large-scale machine learning models, presumably talking substantial steps towards demonstrating the practical utility of quantum computing.
+As such, this work represents one of the most direct and pragmatic applications of quantum algorithms to large-scale machine learning models, presumably taking substantial steps towards demonstrating the practical utility of quantum computing.
 I Introduction
@@
 Unlike traditional models of machine learning, which often rely on explicit likelihood estimation or adversarial training, diffusion models generate data through a gradual, iterative process inspired by the physical concept of diffusion.
-They work by initially introducing noise a data points and then learning how to reverse this process, effectively denoising the data step-by-step.
+They work by initially introducing noise at data points and then learning how to reverse this process, effectively denoising the data step-by-step.
 This framework allows them to generate high-quality samples, often rivaling or surpassing the performance of other generative models, such as generative adversarial networks (GANs) [5] and variational autoencoders (VAEs) [6].
 The appeal of diffusion models lies in their stability during training and their ability to model intricate, high-dimensional data distributions without requiring complex adversarial setups or delicate balancing of training dynamics.
-Diffusion probabilistic models (DPM), as we call them, are families of generative model renowned for their ability to produce high-quality out- puts in complex tasks such as image and audio generation.
+Diffusion probabilistic models (DPM), as we call them, are families of generative models renowned for their ability to produce high-quality out- puts in complex tasks such as image and audio generation.
 Turning to quantum approaches, quantum computing is widely regarded as one of the most promising alternatives to von Neumann architectures in the post-Moore era [7].
@@
 These include factoring [8], database search [9], simulating complex quantum systems [10], sampling [11], and solving problems in linear algebra [12, 13].
-Although we are currently still in what can be called the noisy intermediate-scale quantum (NISQ) era, the concept of a megaquop machine offers a vision of a future where quantum computing achieves remarkable reliability and scalability.
-Such a machine, with an error rate per logical gate on the order of $10^{-6}$, would surpass the limitations of classical, NISQ, or analog quantum devices.
-It would have the capacity to execute quantum circuits involving around $100$ logical qubits and a circuit depth of approximately $10,000$—capabilities far beyond the reach of current technology.
+Although we are currently still in what can be called the noisy intermediate-scale quantum (NISQ) era, the concept of a megaquop machine [14] offers a vision of a future of fault-tolerant application-scale quantum (FASQ) computing [15] in which quantum computing achieves remarkable reliability and scalability.
+Such a megaquop machine, with an error rate per logical gate on the order of $10^{-6}$, would surpass the limitations of classical, NISQ, or analog quantum devices.
+It would have the capacity to execute quantum circuits involving around $100$ logical qubits and a circuit depth of approximately 10,00010,000—capabilities far beyond the reach of current technology.
 Figure 1: Sketch of classical (1.) diffusion probability models (DPM) and the quantum algorithms proposed here for quantum analogs (2.-4.).
-We pursue two approaches: This is on the one hand Carlemann linearization for DPM solvers (2.), on the other Carleman linearization for UniPC (3.).
+We pursue two approaches: This is on the one hand Carleman linearization for DPM solvers (2.), on the other Carleman linearization for UniPC (3.).
 For both approaches, quantum ODE solvers are instrumental in the step before measurement.
@@
 In this section, we lay out our specific results, both concerning rigorous statements that heuristic numerical statements that are meant to provide further evidence for the functioning of the approach.
-Substantial more detail on all steps will be presented in the supplemental material.
+Substantially more detail on all steps will be presented in the supplemental material.
 We will now explain what steps we take towards establishing a fresh direction for demonstrating the utility of quantum computers in machine learning tasks.
@@
 These algorithms underscore the potential of quantum Carleman linearization in addressing diverse mathematical structures.
-Moreover, as we advance toward the era of megaquop machines, this work stands out as one of the most direct and practical applications of quantum algorithms to large-scale machine learning models.
+Moreover, as we advance toward the era of the era of FASQ [15], this work stands out as one of the most direct and practical applications of quantum algorithms to large-scale machine learning models.
 It marks, so we hope, a significant milestone in showcasing the tangible utility of quantum computing in real-world applications.
@@
 In this expression $\lambda\coloneqq\log(\alpha_{t}/\sigma_{t})$ is what is called the log-SNR variable.
-For a given a set of time steps $\{t_{i}\}_{i=0}^{M}$ with $t_{0}=T$ and $t_{M}=0$, with $h_{i}=\lambda_{t_{i}}-\lambda_{t_{i-1}}$, one makes use of the $(k-1)$-st-order Taylor approximation in order to develop what is called the DPM-solver-$k$ in Ref.
+For a given a set of time steps $\{t_{i}\}_{i=0}^{M}$ with $t_{0}=T$ and $t_{M}=0$, with $h_{i}=\lambda_{t_{i}}-\lambda_{t_{i-1}}$, one makes use of the (k1)(k-1)-st-order Taylor approximation in order to develop what is called the DPM-solver-$k$ in Ref.
 [25].
@@
   • Setting $Y=(y_{1},\ldots,y_{N})$, we then derive a linear iteration of the form
-    $\begin{split}\delta\hat{Y}&=\hat{Y}(t_{i})-\hat{Y}(t_{i-1})=A\hat{Y}(t_{i-1})+b\leavevmode\nobreak\ ,\\ &\text{ for }\quad\hat{y}_{1}(t_{M})=\hat{y}_{1}(0)=x_{\mathrm{in}}(0)\leavevmode\nobreak\ ,\end{split}$ (4)
+    $\begin{split}\delta\hat{Y}&=\hat{Y}(t_{i})-\hat{Y}(t_{i-1})=A\hat{Y}(t_{i-1})+b\penalty 10000\ ,\\ &\text{ for }\quad\hat{y}_{1}(t_{M})=\hat{y}_{1}(0)=x_{\mathrm{in}}(0)\penalty 10000\ ,\end{split}$ (4)
     based on Eq. (3), introducing a matrix $A$, which is the quantum Carleman matrix (QCM) associated with the DPM-solver.
@@
   • In terms of $Y=(y_{1},\ldots,y_{N})$, and $Z=(z_{1},\ldots,z_{N})$, we can build on Eq. (5) and Eq. (6) to derive a multi-step linear iteration (predictor) expression given by
-    $\begin{split}\delta Y(t_{i})=\leavevmode\nobreak\ &A^{(0)}Y^{(0)}(t_{i})+b^{(0)}\\ &+A^{(1)}Y^{(1)}(t_{i})+b^{(1)}\\ &+\ldots+A^{(p-1)}Y^{(p-1)}(t_{i})+b^{(p-1)}\leavevmode\nobreak\ ,\end{split}$ (7)
+    $\begin{split}\delta Y(t_{i})=\penalty 10000\ &A^{(0)}Y^{(0)}(t_{i})+b^{(0)}\\ &+A^{(1)}Y^{(1)}(t_{i})+b^{(1)}\\ &+\ldots+A^{(p-1)}Y^{(p-1)}(t_{i})+b^{(p-1)}\penalty 10000\ ,\end{split}$ (7)
     with Carleman matrices $A^{(0)},\dots,A^{(p-1)}$ for all interpolation steps. The corrector, reflecting the other linear iteration takes the form
-    $\begin{split}\delta Z(t_{i})=BZ(t_{i})&+A^{(0)_{c}}Y^{(0)}(t_{i})+b^{(0)_{c}}\\ &+A^{(1)_{c}}Y^{(1)}(t_{i})+b^{(1)_{c}}\\ &+\ldots+A^{(p)_{c}}Y^{(p)}(t_{i})+b^{(p)_{c}}\leavevmode\nobreak\ ,\end{split}$ (8)
+    $\begin{split}\delta Z(t_{i})=BZ(t_{i})&+A^{(0)_{c}}Y^{(0)}(t_{i})+b^{(0)_{c}}\\ &+A^{(1)_{c}}Y^{(1)}(t_{i})+b^{(1)_{c}}\\ &+\ldots+A^{(p)_{c}}Y^{(p)}(t_{i})+b^{(p)_{c}}\penalty 10000\ ,\end{split}$ (8)
     with Carleman matrices $B,A^{(0)_{c}},\dots,A^{(p)_{c}}$.
@@
 A defining feature of the quantum Carleman algorithm for UniPC is the generation of a unique quantum Carleman matrix (QCM) for each interpolation time step $s_{m}$ that we think is also interesting as an innovative step in its own right.
+A practical limitation of the quantum Carleman linearization lies in the truncation of the polynomial hierarchy, which introduces a systematic truncation error that is difficult to control analytically.
+The contribution of the neglected higher-order terms generally depends on the nonlinearity strength and on how strongly probability amplitudes populate higher moments during evolution [29].
+In complex diffusion models, this error may grow over time or under strong coupling, and we do not yet have a rigorous bound for it.
+We emphasize this as an open technical limitation rather than a hidden assumption.
+That said, for many physically relevant regimes—particularly when $N$ is large and the diffusion dynamics remains near a stable manifold—the effect of higher-order monomials is expected to be suppressed.
+Empirically, for smooth drifts and moderate noise levels, low-order truncations reproduce accurate dynamics because higher-order correlations decay rapidly or remain small.
+Thus, although the truncation error is not strictly controllable, it can remain practically negligible within the accuracy relevant for the applications studied here.
+A full characterization of this error scaling and its dependence on truncation order $N$ and diffusion strength, and truncation order is left for future work.
 Figure 2: Results of numerical experiments.
@@
 (4) can be captured in a vector equation
-$M\bm{Y}=\bm{b}\leavevmode\nobreak$ (9)
+$M\bm{Y}=\bm{b}\penalty 10000\$(9)
 that collects $Y(t_{i})$ at all time steps in order to form a single big vector $\bm{Y}=(Y(t_{0}),Y(t_{1}),\dots,Y(t_{M}))$.
 Similarly, the UniPC system can also be captured in a single vector equation as
-$\bm{\mathsf{M}^{c}}\bm{\mathsf{Z}}=\bm{\mathsf{b}^{c}}\leavevmode\nobreak\ ,$ (10)
+$\bm{\mathsf{M}^{c}}\bm{\mathsf{Z}}=\bm{\mathsf{b}^{c}}\penalty 10000\ ,$ (10)
 where $\bm{\mathsf{Z}}$ collects all of the $Z(t_{i})$. The explicit forms of Eq. (9) and Eq. (10) are provided in the supplementary material.
@@
 LCHS achieves gate efficiency by combining linear combinations of unitaries (LCU) [38] with existing Hamiltonian simulation techniques.
-The method involves truncating the infinite integral to a finite interval $[-K,K]$ and discretizing it using the $M+1$ grid points.
+The method involves truncating the infinite integral to a finite interval [K,K][-K,K] and discretizing it using the $M+1$ grid points.
 Compared to the standard QLSS approach, LCHS directly implements the time evolution operator, significantly reducing the number of state preparation oracles required.
@@
 If $M$ is sparse and block-encoded efficiently, the quantum state proportional to the solution vector can be prepared with query complexity scaling as
-$\mathcal{O}(J\kappa\cdot\mathrm{polylog}(N/\epsilon))\leavevmode\nobreak\ ,$ (11)
+$\mathcal{O}(J\kappa\cdot\mathrm{polylog}(N/\epsilon))\penalty 10000\ ,$ (11)
 where $\kappa>0$ is an upper bound on the condition number of $M$. The algorithm achieves a success probability $p_{\text{succ}}$ greater than $1/2$ and accuracy $\epsilon>0$.
@@
 If $\bm{\mathsf{M}^{c}}$ is sparse and efficiently block-encoded, the quantum state proportional to the solution can be prepared with query complexity
-$\mathcal{O}(pJ\kappa\cdot\mathrm{polylog}(N/\epsilon))\leavevmode\nobreak\ ,$ (12)
+$\mathcal{O}(pJ\kappa\cdot\mathrm{polylog}(N/\epsilon))\penalty 10000\ ,$ (12)
 where $\kappa>0$ bounds the condition number of $\bm{\mathsf{M}^{c}}$. The algorithm achieves a success probability $p_{\text{succ}}$ greater than $1/2$ and accuracy $\epsilon>0$.
@@
 II.4 Numerical analysis and experiments
-Above, we present present performance guarantees of the quantum algorithms introduced.
+Above, we present the performance guarantees of the quantum algorithms introduced.
 It is a core contribution of this work, however, to also make the point that the approach taken is reasonably applicable to real-world data.
@@
 where $a_{i}(t)\coloneqq{\lambda_{i}(t)}/{\max_{i,t}|\lambda_{i}(t)|}$.
-By construction, $P(t)$ grows unbounded when negative eigenvalues dominate and decays towards zero when positive eigenvalues prevail.
+By construction, $t\mapsto P(t)$ grows unboundedly when negative eigenvalues dominate and decays towards zero when positive eigenvalues prevail.
 As shown in Fig.
@@
 This work, suggests, however, that there may be significant potential for quantum computers in performing inference tasks or generative tasks of this nature.
+What is more, the present ansatz suggests that the quantum advantage is not only present for highly engineered instances, but in fact robust for an entire class of meaningful instances [15].
 Considering that inference costs might be more than training costs in many machine learning applications, our proposed approach could pave the way for a practical example of quantum advantage in machine learning industry, while also providing a pathway to making quantum computers more useful for solving industrially relevant problems.
@@
 E.
-has been supported by the BMBF (Hybrid++, DAQC, QSolid), the QuantERA (HQCC), the BMWK (EniQmA), the DFG (CRC 183), the Quantum Flagship (PasQuans2, Millenion), the Munich Quantum Valley, Berlin Quantum, and the European Research Council (DebuQC).
+has been supported by the BMFTR (Hybrid++, DAQC, QSolid), the QuantERA (HQCC), the BMWK (EniQmA), the DFG (CRC 183), the Quantum Flagship (PasQuans2, Millenion), the Munich Quantum Valley, Berlin Quantum, and the European Research Council (DebuQC).
 J.-P.